OpenStudy (zeidyt):

is −11π rational or irrational?

OpenStudy (phi):

is pi rational or irrational? any idea ?

OpenStudy (zeidyt):

irrational i believe

OpenStudy (phi):

yes, pi is irrational (you can't write it as a ratio of two numbers) and if you multiply irrational times rational (-11 can be written as the ratio -11/1 so it's rational) you get irrational usually irrational times rational gives irrational. but there are exceptions: \[ \sqrt{2} \cdot \sqrt{2} \] that is irrational times irrational, but the answer is 2 which is rational

OpenStudy (zeidyt):

thank you